OpenStudy (anonymous):

r(t) = t^(6)i-6tj. What is r'(t)?

OpenStudy (amistre64):

the derivative of r ...

OpenStudy (anonymous):

yes, I know that but I keep getting the answer wrong when I'm actually solving for it.

OpenStudy (amistre64):

what do you get when you try to solve?

OpenStudy (amistre64):

the parts are pretty basic, so im not real sure what the difficulty is. If I could see your workings I prolly can determine the error

OpenStudy (anonymous):

I'm not very good at calc so I'm trying to teach myself, sorry. I get 6t^(2)i-6j.

OpenStudy (amistre64):

6-1 is not 2

OpenStudy (anonymous):

so I just replace the 2 with 5 and it's right?

OpenStudy (amistre64):

from what youve posted .. yes r(t) = x(t) , y(t) r'(t) = x'(t) , y'(t) x = t^6 x' = 6t^5 y = -6t y' = -6

OpenStudy (anonymous):

alright thank you. much easier to understand than the explanation in the book.